jay is twice as old as john. 10 years ago,he was 4 times as old as john. Find Jay's present age?

Answer by addingup(3677)

Jay = 2*John We'll use this value for Jay below.
Jay-10 = 4(John-10)
Jay-10 = 4John-40 substitute for Jay per above
(2John)-10 = 4John-40
2John = 4John-30
-2John = -30
John = 15
And since Jay is 2*John:
Jay = 15*2 = 30
----------------------
You can substitute the names for x and y if that's what you're supposed to do